OpenStudy (anonymous):

Whats 1/16 as a power of 1/2

OpenStudy (mathmale):

1 raised to any power is still 1. 2 raised to which power is 16? Then, \[(\frac{ 1 }{ 2 })^n=\frac{ 1 }{ 16 }\] ... it's your job to find n.

jhonyy9 (jhonyy9):

then 2^4 =16 can you finish it ?